P(G) = 0.3 P(G|H) = 0.01 Find the upper or lower bound on P(H).
Answer #1

P(G) = 0.3

P(G|H) = 0.01

P(G and H)/P(H) = 0.01

P(G and H) = 0.01 P(H)

P(G or H) = P(G) + P(H) - P(G and H)

= 0.3 + P(H) - 0.01 P(H)

= 0.3 + 0.99 P(H)

max value of P(G or H) can be 1

hence

1 = 0.3 + 1.01 P(H)

maximum of P(H) = 0.707071

upper bound on P(H) =0.707071

Please rate